Jane found an account at Big Bank that pays out 2.05% compounded monthly. If she needs $1300 in 4 years, how much should she place in the account? Assume that she never adds additional money to the account.
Answer by Theo(13342)

f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

2.05% compounded monthly becomes .0205/12.
the formula uses the rate, not the percent.

4 years * 12 months per year = (4 * 12) months.

formula becomes:

1300 = p * (1 + .0205/12) ^ (4 * 12)

solve for p to get:

p = 1300 / ((1 + .0205/12) ^ (4 * 12)) = 1197.737339.

round to the nearest penny to get 1197.74.

that's how much she needs to invest today at 2.05% per year compounded monthly so that she can have 1300 at the end of the 4 year investment period.